solve for x:
x^2+5x=24

Answer by mananth(16949)

x^2+5x=24
x^2+5x-24=0
x^2+8x-3x-24=0
x(x+8)-3(x+8)=0
(x+8)(x-3)=0
x=-8 OR x=3